User name Huber
Log entry time 15:54:08 on November17,2000
Entry number 52253
keyword=Day shift summary
08:00 - Shift starts with beam still down from CHL failure.
09:20 - Injector Mott measurement now being done. Result is 72.91+/-0.66stat+/-0.48sys
10:38 - good beam is finally re-established.
12:47 - Eugene is here for Moller measurement.
15:27 - Beam back to normal after Moller. Result is -67.82+/-0.11stat+/-2.4%sys
Run Start Stop Events Comments
1372 06:00 09:15 94158 Cosmics run, ended at start of pulsed beam.
1373 10:38 11:45 1M FPP ch1 now looks good.
1374 11:45 12:47 0.97M EOR for Moller.
1375 Run failed because of Bates guys screwing with Right Spect DAQ.
1376 15:30 - continuing at end of shift -